OpenStudy (aonz):

Help please! Differentiate y = x^2 + bx + c, and hence find b and c when x = 3 the gradient is 5, and x = 2 is a zero,

OpenStudy (anonymous):

\[\frac{ dy }{ dx } = 2x + b \] When \[x = 3, \frac{ dy }{ dx } = 5\] So we get \[5 = 6 + b\] \[\rightarrow b = -1\] \[x^2 + bx + c = 0 \rightarrow (2)^2 +(-1)(2) + c = 0 \rightarrow c = -2\]

OpenStudy (aonz):

oh i get how you did it now :D thanks!!